OpenStudy (madeleine2):

Select all the expressions that cannot be factored using integral coefficients and integral constants. a. 25x^2 - 40x + 16 b. 2x^2 + x + 4 c. 6x^2 - 2x - 20 d. x^2 - 2x - 4

OpenStudy (candy13106):

u use guess and check when u factor

OpenStudy (tkhunny):

Do you have he "Discriminant"?

OpenStudy (will.h):

You can do that by finding the GCF

OpenStudy (will.h):

The 1st one doesn't have GCF (greatest common factor) neither for variables nor constants

OpenStudy (tkhunny):

Everything has a GCF.

OpenStudy (will.h):

Integral?

OpenStudy (triciaal):

are they asking which product (AC) will not have factors that add up to B ?

OpenStudy (will.h):

Factoring by groups you mean?

OpenStudy (tkhunny):

Perfect Squares from the Discriminant a. 25x^2 - 40x + 16 ==> 40^2 - 4(25)(16) = 0 ==> You can factor this one. b. 2x^2 + x + 4 ==> 1^1 - 4(2)(4) = -15 ==> Can't factor this one. c. 6x^2 - 2x - 20 ==> 2^2 - 4(6)(-20) = 484 ==> You can factor this one.